OpenStudy (roadjester):

Use the Comparison Theorem to determine whether the integral is convergent or divergent.

OpenStudy (roadjester):

\[\int\limits_{0}^{1}{(\sec x)^{2}\over x \times \sqrt x}dx\]

OpenStudy (anonymous):

i am thinking you can get rid of the secant

OpenStudy (anonymous):

on the interval (0,1), \(sec(x)=\frac{1}{\cos(x)}\) is largest at 0 (where it is 1) and smallest at 1

OpenStudy (anonymous):

so you have \[\int_0^1\frac{1}{\cos^2(x)x^{\frac{3}{2}}}\]

OpenStudy (anonymous):

The integral \[\int_0^1 \frac {dx} { x^{3/2}} \] is divergent.

OpenStudy (anonymous):

think this does not converge, since your problem is at x = 0 and \[\int\frac{dx}{x^{\frac{3}{2}}}\] is divergent